OpenStudy (anonymous):

let z and w be complex numbers such that z+w=0 and z square+w square=1,then mod of z-w=?.

OpenStudy (perl):

so you want the modular distance of z-w

OpenStudy (perl):

| z-w| = sqrt ( z-w)^2 =

OpenStudy (perl):

, sorry messed up

OpenStudy (perl):

| z- w | = (z-w) conjugat ( z- w)

OpenStudy (perl):

you need some formulas http://en.wikipedia.org/wiki/Complex_conjugation

OpenStudy (perl):

|z|^2 = z* z' , where z' is the conjugate

OpenStudy (perl):

and |z| is mod z

OpenStudy (anonymous):

thanks fr ur help but i was not getting the answer i got it now :)

OpenStudy (perl):

can you type out the answer :)

OpenStudy (perl):

z^2 + w^2 = 1 (z + wi) (z-wi)= 1

OpenStudy (anonymous):

2

OpenStudy (anonymous):

do u want a solution?

OpenStudy (perl):

yes

OpenStudy (anonymous):

mark the two equations as (1) and (2) from (1),we get w=-z, then 2z square=1 z=+_1/root2 for z=+1/root2,w=-1/root2 for z=-1/root2,w=+1/root2 so,mod of z-w=root2

OpenStudy (anonymous):

sorry,i forgot to write root in the previous answr hope you ll get this solution its written in a bit puzzling manner :P

OpenStudy (perl):

makes sense, mod of z-w just becomes the real number root 2

OpenStudy (anonymous):

hmmm..right !